OpenStudy (anonymous):

If arg(z)<0 , then arg(-z) - arg(z) equal to ?

OpenStudy (anonymous):

@satellite73 @mukushla @eliassaab @sami-21 @sauravshakya

OpenStudy (anonymous):

\[z=x+iy\]then\[\arg(z)=\arctan \frac{y}{x}\] ?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

-z = -x - iy arctan = -y/-x

OpenStudy (anonymous):

y/x - y/x = 0 so......tan 180

OpenStudy (anonymous):

Yup....thxx.....